

********************************************************************
UARS WEEKLY STATUS REPORT - 30 April 2004
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to perform nominally, taking solar data on all
available orbital tracking periods. This week we received about 65% of
each day's data amounting to nearly 605 minutes of solar observing time
per day.
***********
SUSIM
***********
SUSIM successfully gathered its daily mid (1.1 nm) resolution, Mg
II, occultation, solar pointing offset, and high resolution
continuous monitoring scans for periods of UARS real-time telemetry
or onboard tape recording.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact for UARS
instrument operations in support of the Science Traceability
Mission of UARS.
Daily processing of instrument Daily Activity Plans continues
with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 17 April 2004 (Orbit 68939, GMT Day 108) through
23 April 2004 (Orbit 69043, GMT Day 114).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 112/1331 69007 PEM Z/H TO FULLTIME
PEM 113/2204 69028 PEM ZEPS/HEPS TO DAY ONLY
HALOE 114/1438 69038 POWER ON
The beta angle ranged between the angles of 68.4 degrees to 53.8
degrees for this report period. The beta angle is now decreasing
towards a minimum angle of 0 degrees on 08 May 2004 (DOY 128).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.6 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 1.26 to 1.27 degrees C. Additional
battery performance data for each battery on 16 April and 23 April
2004 are provided in Appendix D.
The clock error ranged from 10.1 to -10.6 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT
for this report. A summary is listed below and the details of ALERT
and Out-Of-Limits occurrences are listed in Appendix B.
OUT-OF-LIMITS = 03
ALERTS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
FOT = 05
PEM = 02
HALOE = 02
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
FOT = 03
Two (2) Anomaly Reports remained OPEN at the end of this report period
(see Appendix E).
GROUND SYSTEM OPERATION
Three (3) operational support problems occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF).
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT was informed on 13 April 2004, that TDRS-8 will be moved to
TDW and TDW will be moved to 172. This transition took place on 23
April 2004 as scheduled. TDRS-8 is a newer spacecraft and uses SMA
service, instead of MA. This service has never been tested or used by
the FOT, so the FOT had to find a temporary solution until this
service can be utilized. The FOT decided to use the new TDW, but only
with the SSA service. Since SSA time is in higher demand the FOT will
lose approximately half of its TDW time. In the meantime, the
software fix to the CMS is being reviewed to ingest the SMA view
periods correctly.
UARS flew through a solar eclipse on 19 April 2004. The eclipse
affected only one orbit and caused a false night indication onboard.
The false night indication caused PEM's day-only operations to cycle
twice during the day instead of just once. Also, the power
calculations onboard were skewed for one orbit. All other systems
remained nominal.
At the meeting with FDF for 20 April 2004, it was determined that the
initial 200-300 arcsecond error reported to us was due to a truncation
issue with the solar ephemeris telemetry points gathered. The actual
error was closer to 20-30 arcseconds. That means that the solar
ephemeris is not the culprit behind the attitude error seen by the
HRDI team. The FOT will continue with its investigation to determine
the real cause.
MEETINGS
The FOT held a meeting with FDF on 20 April 2002 to discuss the
onboard solar ephemeris issue. At this meeting it was found that the
initial calculation of 200-300 arcsecond error was in error and it was
actually 20-30 arcseconds. The error was induced by truncation of
telemetry and limitations of the onboard computer's algorithms.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 07 May 2004 (DOY 128). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 23 April 2004
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
68946 108/1123 04-115 OBC/CLK CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
68671 110/0249 04-116 OBC/CLK CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
68996 111/1847 04-117 OBC OBC MEM MON
68999 112/0026 04-118 OBC/CLK CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
69007 112/1331 04-119 PEM PEM Z/H TO FULLTIME
69022 113/1224 04-120 OBC/CLK CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
69028 113/2204 04-121 PEM PEM ZEPS/HEPS TO DAY ONLY
69038 114/1438 04-122 HALOE POWER ON
69039 114/1620 04-123 HALOE POWER ON CONTINUE
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 23 April 2004
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
PEA1ONNOARM R-LO AXIS 1 BIAS ON ERR 68753
PEA2LV7E R-LO AXIS 2 -6V 68753
PEA2LV4E R-LO AXIS 2 -14V 68753
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 23 April 2004
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
68946 108/1123 CLK ADJ FR 20.6 TO 41.2 MSEC/DAY
68971 110/0254 CLK ADJ FR 41.2 TO 20.6 MSEC/DAY
68999 112/0026 CLK ADJ FR 20.6 TO 0.0 MSEC/DAY
69022 113/1224 CLK ADJ FR 0.0 TO 20.6 MSEC/DAY
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 23 April 2004
16 April 2004 - GMT Day 107 Beta = 67.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.31 +00.0/+00.0 07.0 0.00 0.0
2 +11.2/-112.0 +3.87/+2.60 +18.8/-14.4 26.4 1.05 15.5
3 +00.0/-050.4 +2.91/+1.65 +18.4/-14.4 26.4 1.05 15.1
23 April 2004 - GMT Day 114 Beta = 53.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 -700.0/-700.0 -3.05/-4.94 +00.0/+00.0 07.0 0.00 0.0
2 +00.0/-089.6 +3.87/+2.60 +22.8/-14.4 25.8 1.05 15.5
3 +00.0/-123.2 +2.91/+1.65 +22.4/-13.6 25.8 1.05 15.1
V/T LEVEL CHANGES
ORB# AOS TIME LVL FR-TO
----- -------- ---------
-None this report period.
APPENDIX E
UARS ANOMALY REPORTS
Week ending 23 April 2004
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-019 FOT 68959 109/0838 30832 Generic Late Acquisition
04-019 FOT 68971 110/0249 30832 Generic Late Acquisition
04-020 FOT 68977 110/1337 N/A Eclipse False Night Indicator
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME CDS PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
04-019 FOT See List Apr 04 30832 Generic Late Acquisition
04-020 FOT 68977 110/1337 N/A Eclipse False Night Indicator